OpenStudy (anonymous):

Let f(x) = x2 - 1 and g(x) = 2x - 3. Find the new function and write it in simplest form. f + g a. -x2 + 2x - 2 c. x2 + 2x - 4 b. x2 - 2x + 2 d. 2x3 - 3x2 - 2x + 3

OpenStudy (anonymous):

c

OpenStudy (math&ing001):

(f+g)(x) = f(x) + g(x) = x2 - 1 + 2x - 3 = x2 + 2x - 4 Welcome to Openstudy !